A 3.0-kg block moving to the right on a frictionless table at 4.0 m/s makes a head-on collision with another 3-kg block moving 5 m/s to the left. Find the final velocity of the blocks, if the collision is PERFECTLY INELASTIC.

1
Expert's answer
2022-01-12T08:31:35-0500

"m_1 = 3 \; kg \\
m_2 = 3 \; kg \\
u = 4 \; m/s \\
v = 5 \; m/s"

By momentum conservation:

"3 \times 4 -3 \times 5 = (3+3)v_f \\
v_f = -0.5 \; m/s"